Of a sample of 850 people, 415 of them have negative tests. 115 people do have Herpes and have a negative test. 250 people do have Herpes and have a positive test.

Based on the information above select the option that has the correct value and interpretation for the Negative Predictive Value of this test.

.Group of answer choices:

.Someone who has a negative test has a 72% chance of not having Herpes.

.Someone who has does not have Herpes has a 72% chance of getting a negative test result.

.Someone who has a negative test has a 62% chance of not having Herpes.

.Someone who has a positive test has a 72% chance of having Herpes.

.Someone who has Herpes has a 62% chance of getting a positive test result.

.Someone who has does not have Herpes has a 62% chance of getting a negative test result.
Answer #1

Answer)

Negative predicted value = (true negative)/(true negative + false negative)

Given info

Total = 850

Negative test = 415

115 people do have herpes and have negative test

That is true negative = 115

False negative = 415 - 115 = 300

So negative predicted value = 115/415 = 28%

Which means someone having this test do not having the herpes would have a negative test 28% of the time

Or Someone who has a negative test has a 72% chance of not having Herpes.